# thanks: These authors contributed equallythanks: These authors contributed equally

### v2

@@ Abstract
Metamaterials are a promising platform for a range of applications, from shock absorption to mechanical computing. These functionalities typically rely on floppy modes or mechanically frustrated loops, both of which are difficult to design. We introduce a combinatorial approach that allows to create an arbitrarily large number of floppy modes and frustrated loops. The design freedom of the mode shapes enables us to easily introduce kinematic incompatibility to turn them into frustrated loops. We demonstrate that floppy modes can be sequentially buckled by using a specific instance of elastoplastic buckling, and we utilize our combinatorial floppy chains and frustrated loops to achieve matrix-vector multiplication in materia. Our findings bring about new principles for the design and the use of floppiness and geometric frustration in soft matter and metamaterials.
@@
Introduction – In soft matter, mechanical properties are often rooted in floppy modes and geometrically frustrated states of self-stress. Examples range from protein allostery [1], granular packings [2, 3] and colloidal glasses [4, 5] to amorphous solids [6, 7, 8], biopolymer networks [9, 10, 11] and mechanical metamaterials [12, 13]. Mechanical metamaterials are a platform of choice to study floppy and frustrated modes [14]. Once one understands those modes, one can in turn harness them to achieve on-demand unusual properties [15]. Prime examples are topological wave guiding and stress focusing [16, 17, 18, 19, 20, 21, 22, 23], vibration and shock absorption [24, 25, 26], shape morphing [27, 28, 29, 30] and mechanical computing [31, 32, 33, 34, 35, 36].
@@
Floppy modes are deformations that cost negligible elastic energy [37, 38, 39, 40, 41, 42, 43, 44]. Frustrated stressed states are often configured in loops [45, 46, 47, 48, 49, 50, 51], where these low-energy deformations are geometrically impossible. Both floppy modes and frustrated loops can be used to channel deformations and stresses efficiently, yet designing metamaterials with multiple, precisely controlled modes or loops remains a challenge. To address this problem, we open up the design space of combinatorial metamaterials, by introducing a theoretical approach that allows much freedom in setting the number and shapes of floppy modes and frustrated loops. We leverage this approach to create metamaterials that exhibit advanced sequential response upon uniaxial compression and mechanical computing in the form of matrix-vector multiplication. Thus, our work bridges abstract principles and practical applications in soft programmable materials and adaptive architectures.
@@
Figure 1: Floppy modes in metamaterials with perfect hinges and rigid bonds. a, Triangular blocks with spins (green) describing the direction of displacement of edge nodes, and bonds (brown) constraining spins to move in alternating manner. One internal bond in Block $T_{1}$ constrains two spins, thus the third is independent, resulting in two floppy modes. Block $T_{2}$ has two internal bonds, thus all spins displace together in one floppy mode. b, Metamaterial made of $T_{1}$ and $T_{2}$ blocks. Chains of connected spins constraining each other’s motion are individually colored; open chains (light green, yellow, orange) represent nodes moving together in a floppy mode, and so do chains that contain closed loops of even length (dark green). Chains with odd loop (red) are mechanically frustrated and hence rigid. c, Experimental demonstration of the four floppy modes of the system in b. d, Normalized number $F/N$ of floppy modes vs. normalized number $N_{1}/N$ of $T_{1}$ blocks for a lattice of $N=210$ blocks. The values of $F$ span almost the entire range between the lower and upper theoretical bounds. The average number of floppy modes in randomly generated systems exhibits small fluctuations and is closer to the lower bound.
@@
Building Blocks and Spin Model – We consider two-dimensional mechanical metamaterials composed of triangular building blocks, which are constructed from rigid bonds connected at freely rotating hinges. Each block consists of three corner nodes and three edge nodes. The nodes are connected by six rigid perimeter bonds. In addition, $T_{1}$ triangles have one internal bond connecting two adjacent edge nodes, while $T_{2}$ triangles have two such internal bonds (Fig. 1a). For floppy modes, in the small deformation limit, the corner nodes remain stationary while edge nodes move perpendicular to their respective edges. This key insight allows us to introduce a spin model to describe the system’s deformations. In this model, we assign a spin-like variable to each edge node, describing whether it displaces in to the triangle or out of it. Effective antiferromagnetic interactions between spins are introduced by internal bonds, since they constrain connected edge nodes to move in alternating directions with respect to the triangle.
@@
Design of Floppy Modes – Using our spin model, we can design metamaterials with varying numbers and shapes of floppy modes. By connecting spins, we create chains of nodes (Fig. 1b) that can move together, independently from other chains in the system (Fig. 1c). The mutual direction of motion of the nodes is determined by the internal bonds connecting them along the chain. This demonstrates our ability to create chains with complex, predefined shapes, and deformations. Furthermore, an inherent property caused due to the antiferromagnetic interaction between spins is that a closed loop with an even number of such bonds allows the corresponding spins to move together (Fig. 1b,c, dark green chain), while a loop with an odd length renders the chain frustrated and rigidifies the connected spins (Fig. 1b, red chain).
@@
Number of Modes – On top of designing the spatial form of floppy modes, the spin model allows to quantify their number in such metamaterials; The number of floppy modes is the number of independent degrees of freedom in the system. Each spin is a degree of freedom, and each bond is a constraint that reduces the number of independent degrees of freedom. However, a bond which closes a loop is redundant and does not add a constraint, while chains with an odd loop are rigid and do not possess a degree of freedom. A lattice of $N$ triangles and of perimeter $P$ contains $\frac{3N}{2}+\frac{P}{2}$ edge nodes, or spins. Denoting the numbers of $T_{1}$ and $T_{2}$ triangles by $N_{1}$ and $N_{2}$, respectively, the number of internal bonds is $N_{1}+2N_{2}$. Using $N=N_{1}+N_{2}$, we obtain an exact expression for the number of floppy modes, $F=N_{1}-\frac{N}{2}+\frac{P}{2}+L-R$, where $L$ is the number of closed loops of internal bonds, and $R$ is the number of rigid chains.
For given lattice size and shape that set $N$ and $P$, and for given values $N_{1}$ and $N_{2}$ of $T_{1}$ and $T_{2}$ blocks, the number of floppy modes can take a wide range (Fig. 1d), depending on how the blocks are arranged to form chains and loops, thus changing the value of $L-R$. The metamaterial design can lead to any number of modes between the bounds,
@@
If most triangles are $T_{2}$ $\left({\rm namely,}\ N_{1}\lesssim\frac{N}{2}\right)$, the triangles can be arranged such that the metamaterial will be rigid, without any floppy modes, $F=0$. For any ratio of $T_{1}$ and $T_{2}$ triangles, there are always designs that give an extensive number of floppy modes $F\propto N$, and if most blocks are $T_{1}$ $\left(N_{1}\gtrsim\frac{N}{2}\right)$, this number is necessarily extensive. Finally, we observe that by randomly positioning and orienting the blocks, the number of floppy modes is closer to the lower bound (Fig. 1d).
@@
Connecting specific edge nodes, we can design floppy modes of arbitrary topology, like catenated floppy loops (Fig. 2c) that may be individually actuated since they do not touch each other (Fig. 2d and Video 2). Furthermore, chains that in individual layers are floppy may now form rigid loops that traverse multiple layers in a manner that is sensitive to the positions of the vertical connectors. For instance, moving the black connector in Fig. 2c to its dotted position changes the parity of the yellow loop from even and thus floppy to odd and thus rigid, without changing the structure of any of the layers.
@@
By extending our design methodology to layered lattices, we significantly expand the range of achievable functionalities and topologies in our mechanical metamaterials and (literally) bridge the gap between planar designs and fully three-dimensional structures. In the remainder of the paper, we utilize (i) floppy modes to achieve functional responses such as sequential buckling of floppy modes with designer shapes and (ii) frustrated loops to realize matrix-vector multiplication.
Figure 3: Sequential actuation of floppy modes. a, Theoretical model and corresponding metamaterial geometry, with rigid bonds replaced by diamonds. b, Metamaterial with three floppy chains (red) of varying length, separated by disconnected rigid chains (blue). c, 3D-printed metamaterial. d, Elastic metamaterial at different stages of compression. f Compression of plastic metamaterial of the same design. h, Compression of plastic metamaterial with three identical floppy chains. Color-coded arrows in d, f, and h indicate horizontal displacement. e, g, i, Corresponding vertical force, $F$ and average horizontal displacement, $u_{1}$ of each floppy chain vs. global compressing stroke $u/H$. $B$, $M$, and $T$ denote the identically shaped bottom, middle, and top floppy chains in h. The sign of the slope ($S$) of force vs compression is detected (dashed lines).
@@
Floppy Mode Buckling – To go beyond local actuation of individual modes, we exploit the buckling instability as a mechanism to actuate floppy modes by uniform global compression. Unlike existing metamaterials based on buckling, with one mode [52], a large number of modes along straight lines [13, 39, 41, 26, 53], or limited modes in a hierarchical design [54, 55], our modes can be shaped into tortuous chains or loops, and their buckling onset is determined by the length of the chains and by the parity of the loops (Video 3).
We demonstrate floppy mode actuation in a 3D-printed metamaterial made of diamond-shaped beams (Fig. 3a); The thin connectors at the diamond ends allow a relatively flexible rotation compared to the substantially greater resistance to deforming the individual diamonds. We design three roughly horizontal floppy chains of different lengths and separate them by rigid layers (Fig. 3b). To trigger the buckling of the floppy modes, we apply a uniform vertical compression to the metamaterial. We study elastic buckling by 3D printing with a rubber-like polymer (Fig. 3c). Under compression, the three floppy modes buckle almost simultaneously (Fig. 3d,e) because of their close buckling loads and the positive stiffness at the onset of elastic buckling (Fig. 3e).
@@
Sequential Yield Buckling – We separate the excitation of the different modes by exploiting yield buckling [26], which is rooted in the symbiotic occurrence of plastic yielding and buckling. To combine plasticity with the floppy modes, we 3D print an elastoplastic metamaterial of the same geometry. Under quasistatic compression, the straight mode of lowest energy buckles first, followed by a negative stiffness, until all the diamonds in the first buckled layer reach self-contact (Fig. 3f left). At this point, the metamaterial stiffens until the critical buckling load of the second-length mode is reached (Fig. 3f middle). Such sequential buckling repeats until the last mode with the longest length buckles and reaches contact. Due to this sequential buckling, the force curve exhibits a wiggly increasing plateau with tunable local maxima (Fig. 3g). Thus, our designed floppy modes can be used to tailor the system’s nonlinear force-displacement response. A metamaterial with all floppy chains of the same length also exhibits sequential buckling with negative stiffness at the onset of yield buckling (Fig. 3h,i). Here, the buckling order of the modes is determined by imperfections and boundary conditions.
@@
Matrix-Vector Multiplication – Computing in materia is an emerging direction [56, 32], studied in a myriad of physical platforms, such as cross-bar arrays [57], spintronic devices [58], electromagnetic [59, 60] and phononic coupled resonators [61], and microfluidics [62]. For these, the ability to perform algebraic operations is paramount, and therefore a crucial challenge is how to achieve matrix manipulations. Recent work has shown that floppy modes can be used to achieve matrix-vector multiplication in mechanical systems [35]. Here, we demonstrate that frustrated loops are an efficient way to perform algebraic operations with multiple inputs and outputs.
@@
Figure 4: Matrix-vector multiplication. a, Experimental setup wherein the six corner nodes are fixed in place. Arrows indicate the decay in displacement along the two branches of the frustrated loop (dashed). b, Experimental results (dots) and theoretical predictions with $\alpha=0.8$ (solid line) of output under a single input from the top, $I_{\text{T}}$ or the bottom, $I_{\text{B}}$. Colors correspond to the outputs marked in panel a. c, Experimental measurements (solid lines) and predictions (dashed lines) under two simultaneous inputs. d, Three-layer metamaterial. e, Corresponding 3D structure. f, Experiment results (solid lines) of inputs in the middle layer, and outputs in the left and right layers. The dashed lines are the predictions.
The minimal metamaterial with a frustrated loop consists of six $T_{2}$ blocks (Fig. 4a). To enable manual actuation, we 3D print the diamond-shaped beams in this metamaterial with a flexible material and print the ligaments with a softer material. We fix the hexagon’s six corner nodes and locally actuate the top and bottom edge nodes, as input displacements $I_{\rm T}$ and $I_{\rm B}$. In the ideal model of rigid bonds, realized above using LEGO®, a chain with an odd loop is rigid, and may not deform. However, the 3D-printed metamaterial has some flexibility, leading to non-trivial output displacements $O_{\rm TL}$, $O_{\rm TR}$, $O_{\rm BL}$, and $O_{\rm BR}$ (Video 4). Actuating each of the inputs separately, we observe linear behavior of each of the outputs (Fig. 4b). Thus, this frustrated loop relates the input, $\vec{I}=\begin{pmatrix}I_{\rm T},I_{\rm B}\end{pmatrix}^{T}$, and output $\vec{O}=\begin{pmatrix}O_{\rm TL},O_{\rm TR},O_{\rm BL},O_{\rm BR}\end{pmatrix}^{T}$ vectors via the matrix-vector product, $\vec{O}=A\vec{I}$.
Each internal bond’s effective antiferromagnetic interaction flips the displacement’s direction. Due to the odd length of frustrated loops, the transmission of displacement from the input node along the loop’s two branches leads to displacement in opposite directions at the output node. However, beyond these flips in direction, the elastic nature of the 3D-printed metamaterial causes a decay in the displacement’s magnitudes [63], which we theoretically describe by each beam decreasing the displacement by some factor $\alpha$. Tracking the two paths from each input to each output node in the hexagonal structure of Fig. 4a, we obtain the transmission matrix,
$A$ $=\begin{pmatrix}-\alpha^{4}+\alpha^{7}&-\alpha^{2}+\alpha^{7}&\alpha^{4}-\alpha^{7}&\alpha^{4}-\alpha^{7}\\ -\alpha^{3}+\alpha^{8}&\alpha^{4}-\alpha^{5}&\alpha^{3}-\alpha^{8}&-\alpha^{4}+\alpha^{7}\end{pmatrix},$ (2)
which is consistent with the experimental measurement (Fig. 4b). Simultaneously applying two inputs in the same direction and then in opposite directions, we see that superposition holds (Fig. 4c), such that the output displacements are indeed given by $\vec{O}=A\vec{I}$.
Each frustrated loop may receive two inputs. Multiplying larger vectors is possible by feeding the output from one loop as an input to another loop. However, this requires transmission lines to cross, which is problematic in 2D metamaterials. We overcome this by coupling adjacent layers in a 3D metamaterial (Fig. 4d,e). Individually actuating each of the input nodes in the middle layer and measuring the response of each of the output nodes in the left and right layers, we obtain all elements of the transmission matrix $A$. We test our approach by simultaneously applying different combinations of several inputs (Fig. 4f), and observe that the output displacements are well described by $\vec{O}=A\vec{I}$.
@@
We introduced a combinatorial approach using unimodal and bimodal triangular building blocks to create metamaterials with arbitrary shapes of multiple floppy modes and frustrated loops. We demonstrated that curved floppy modes together with yield buckling can be used to achieve sequential deformations, and that frustrated loops can be exploited to achieve matrix-vector multiplication. Our findings enrich the toolbox of sequential metamaterial design and lead to various applications in shock and vibration damping, shape changing, and mechanical computing. Our work opens questions like: can we use plasticity and fatigue to learn in combinatorial metamaterials [64]? How can frustrated loops integrate matrix-vector multiplication and nonlinear response like buckling for the function of logic gates [49]? How can floppy modes and frustrated loops be used for locomotion in active metamaterials [65]?
@@
Acknowledgments – We thank Chaviva Sirote-Katz, Daniela Kraft, Dor Shohat, Julio Melio and Marc Serra-Garcia for inspiring discussions and Bat-El Pinchasik and Yaara Shokef for technical assistance. T.A.S. and Y.S. thank the Institute of Physics at the University of Amsterdam for its hospitality. T.A.S. acknowledges funding from the Israeli Ministry of Energy and Infrastructure’s scholarship program for undergraduate students in the field of energy. W.L. and C.C. acknowledge funding from the European Research Council under grant agreement 852587 and the Netherlands Organisation for Scientific Research under grant agreement NWO TTW 17883. C.C. acknowledges funding from the Netherlands Organisation for Scientific Research under grant agreement NWO under grant agreement VI.Vidi 2131313. This research was supported in part by the Israel Science Foundation Grant No. 1899/20.
@@
Code and Data Availability – The supplementary videos as well as all the data and computer codes supporting this work are available at https://doi.org/10.5281/zenodo.15130951.

### v3

# Combinatorial Design of Floppy Modes and Frustrated Loops in Metamaterials

@@ Abstract
Metamaterials are a promising platform for a range of applications, from shock absorption to mechanical computing. These functionalities typically rely on floppy modes or mechanically frustrated loops, both of which are difficult to design. In particular, how to design multiple modes or loops with target deformations remains an open problem. We introduce a combinatorial approach that allows us to create an arbitrarily large number of floppy modes and frustrated loops. The design freedom of the mode shapes enables us to easily introduce kinematic incompatibility to turn them into frustrated loops. We demonstrate that floppy modes can be sequentially buckled by using a specific instance of elastoplastic buckling. We utilize our combinatorial floppy chains and frustrated loops to achieve matrix-vector multiplication in materia. Our findings bring about new principles for the design and the use of floppiness and geometric frustration in soft matter and metamaterials.
@@
Introduction – In soft matter, mechanical properties are often rooted in floppy modes and geometrically frustrated states of self-stress. Examples range from protein allostery [1], granular packings [2, 3] and colloidal glasses [4, 5] to amorphous solids [6, 7, 8], biopolymer networks [9, 10, 11] and mechanical metamaterials [12, 13, 14]. Mechanical metamaterials are a platform of choice to study floppy and frustrated modes [15]. Once one understands those modes, one can in turn harness them to achieve on-demand unusual properties [16]. Prime examples are topological wave guiding and stress focusing [17, 18, 19, 20, 21, 22, 23], vibration and shock absorption [24, 25, 26], control of fracture [27, 28, 29], shape morphing [30, 31, 32, 33] and mechanical computing [34, 35, 36, 37, 38, 39].
@@
Floppy modes are deformations that cost negligible elastic energy [40, 41, 42, 43, 44, 45, 46, 47]. Frustrated stressed states are often configured in loops [48, 49, 50, 51, 52, 53, 54], where these low-energy deformations are geometrically impossible. Both floppy modes and frustrated loops can be used to channel deformations and stresses efficiently. Yet, designing metamaterials with multiple, precisely controlled modes or loops remains a challenge. Existing works rely on topological [12, 13, 17, 11, 55, 14] or computational [1, 44, 46] approaches and exhibit limited design freedom on the number of modes and loops, and on their spatial structure.
Here, we open up the design space of combinatorial metamaterials, by introducing a theoretical strategy that allows unpredecented freedom in setting the number and shapes of floppy modes and frustrated loops. We leverage this approach to create metamaterials that exhibit advanced sequential response upon uniaxial compression and mechanical computing in the form of matrix-vector multiplication. Thus, our work bridges abstract principles and practical applications in soft programmable materials and adaptive architectures.
@@
Figure 1: Floppy modes in metamaterials with perfect hinges and rigid bonds. a, Triangular blocks with spins (blue arrows) describing the direction of displacement of edge nodes, and bonds (brown) constraining spins to move in alternating manner. One internal bond in Block $T_{1}$ constrains two spins, thus the third is independent, resulting in two floppy modes. Block $T_{2}$ has two internal bonds, thus all spins displace together in one floppy mode. b, Metamaterial made of $T_{1}$ and $T_{2}$ blocks. Chains of connected spins constraining each other’s motion are individually colored; open chains (light green, yellow, orange) represent nodes moving together in a floppy mode, and so do chains that contain closed loops of even length (dark green). Chains with odd loop (red) are mechanically frustrated and hence rigid. c, Experimental demonstration of the four floppy modes of the system in b. d, Normalized number $F/N$ of floppy modes vs. normalized number $N_{1}/N$ of $T_{1}$ blocks for a lattice of $N=210$ blocks. The values of $F$ span almost the entire range between the lower and upper theoretical bounds. The average number of floppy modes in randomly generated systems exhibits small fluctuations and is closer to the lower bound.
@@
Building Blocks and Spin Model – We consider two-dimensional mechanical metamaterials composed of triangular building blocks, which are constructed from rigid bonds connected at freely rotating hinges. Each block consists of three corner nodes and three edge nodes. The nodes are connected by six rigid perimeter bonds. In addition, $T_{1}$ triangles have one internal bond connecting two adjacent edge nodes, while $T_{2}$ triangles have two such internal bonds (Fig. 1a). For floppy modes, in the small deformation limit, the corner nodes remain stationary while edge nodes move perpendicular to their respective edges. This key insight allows us to introduce a spin model to describe the system’s deformations. In this model, we assign a binary spin-like variable to each edge node, describing whether it displaces into the triangle or out of it. Effective antiferromagnetic interactions between spins are introduced by internal bonds, since they constrain connected edge nodes to move in alternating directions with respect to the triangle. Thus, a floppy, or zero-energy deformation corresponds to a spin configuration which satisfies all the antiferromagnetic interactions, while frustration appears whenever there is a closed loop with an odd number of bonds.
@@
Design of Floppy Modes – Using our spin model, we can design metamaterials with varying numbers and shapes of floppy modes. By connecting spins, we create chains of nodes (Fig. 1b and Video 1) that can move together, independently from other chains in the system (Fig. 1c). The mutual direction of motion of the nodes is determined by the internal bonds connecting them along the chain. This demonstrates our ability to create chains with complex, predefined shapes, and deformations. Furthermore, an inherent property caused due to the antiferromagnetic interaction between spins is that a closed loop with an even number of such bonds allows the corresponding spins to move together (Fig. 1b,c, dark green chain), while a loop with an odd length renders the chain frustrated and rigidifies the connected spins (Fig. 1b, red chain).
@@
Number of Modes – On top of designing the spatial form of floppy modes, the spin model allows to quantify their number in such metamaterials; the number of floppy modes is the number of independent degrees of freedom in the system. Each spin is a degree of freedom, and each bond is a constraint that reduces the number of independent degrees of freedom. However, a bond which closes a loop is redundant and does not add a constraint, while chains with an odd loop are rigid and do not possess a degree of freedom. A lattice of $N$ triangles and of perimeter $P$ contains $\frac{3N}{2}+\frac{P}{2}$ edge nodes, or spins. Denoting the numbers of $T_{1}$ and $T_{2}$ triangles by $N_{1}$ and $N_{2}$, respectively, the number of internal bonds is $N_{1}+2N_{2}$. Using $N=N_{1}+N_{2}$, this analysis yields an exact expression for the number of floppy modes, $F=N_{1}-\frac{N}{2}+\frac{P}{2}+L-R$, where $L$ is the number of closed loops of internal bonds, and $R$ is the number of rigid chains.
For given lattice size and shape that set $N$ and $P$, and for given values $N_{1}$ and $N_{2}$ of $T_{1}$ and $T_{2}$ blocks, the number of floppy modes can take a wide range (Fig. 1d), depending on how the blocks are arranged to form chains and loops, thus changing the value of LRL-R. The metamaterial design can lead to any number of modes between the bounds (see Appendix),
@@
If most triangles are $T_{2}$ $\left({\rm namely,}\ N_{1}\lesssim\frac{N}{2}\right)$, the triangles can be arranged such that the metamaterial will be rigid, without any floppy modes, $F=0$. For any ratio of $T_{1}$ and $T_{2}$ triangles, there are always designs that give an extensive number of floppy modes $F\propto N$, and if most blocks are $T_{1}$ $\left(N_{1}\gtrsim\frac{N}{2}\right)$, this number is necessarily extensive. Finally, we observe that by randomly positioning and orienting the blocks, the number of floppy modes is closer to the lower bound (Fig. 1d, and see Appendix). In summary, our combinatorial strategy allows us to design an arbitrary number of modes in a systematic way and with complex target shapes, unlike existing approaches thus far [12, 17, 13, 1, 11, 55, 44, 46, 14].
@@
Connecting specific edge nodes, we can design floppy modes of arbitrary topology, like catenated floppy loops (Fig. 2c) that may be individually actuated since they do not touch each other (Fig. 2d and Video 2). Furthermore, chains that in individual layers are floppy may now form rigid loops that traverse multiple layers in a manner that is sensitive to the positions of the vertical connectors. For instance, moving the black connector in Fig. 2c to its dotted position changes the parity of the yellow loop from even and thus floppy to odd and thus rigid, without changing the structure of any of the layers. Generalizing the analysis of the number of floppy modes to such layered metamaterials is straightforward (see Appendix).
@@
By extending our design methodology to layered lattices, we significantly expand the range of achievable functionalities and topologies in our mechanical metamaterials and (literally) bridge the gap between planar designs and fully three-dimensional structures.
In the remainder of the paper, we utilize floppy modes and frustrated loops to achieve functional responses such as (i) the sequential buckling of designer-shaped floppy modes under uniform compression and (ii) matrix-vector multiplication with multiple inputs and outputs under local actuation. In both, the enhanced design freedom of our combinatorial approach leads to more advanced responses than in earlier works.
@@
Floppy Mode Buckling – To go beyond local actuation of individual modes, we exploit the buckling instability as a mechanism to actuate floppy modes by uniform global compression. Unlike existing metamaterials based on buckling, with one mode [56], a large number of modes along straight lines [13, 42, 44, 26, 57], or limited modes in a hierarchical design [58, 59], our modes can be shaped into tortuous chains or loops based on the combinatorial design of the metamaterial, and their buckling onset is determined by the length of the chains and by the parity of the loops (Video 3).
We demonstrate floppy mode actuation in a 3D-printed metamaterial made of diamond-shaped beams (Fig. 3a). The thin connectors at the diamond ends allow a relatively flexible rotation compared to the substantially greater resistance to deforming the individual diamonds. We first demonstrate two identical floppy chains separated by rigid regions in an irregular-shaped metamaterial with two isolated holes inside (Fig. 3b)
Figure 3: Sequential actuation of floppy modes. a, Theoretical model, corresponding metamaterial geometry and 3D printed components, with rigid bonds replaced by diamonds. b, Irregularly shaped metamaterial with two identical floppy chains (red), separated by disconnected rigid chains (blue). e, Metamaterial with three floppy chains (red) of varying length, separated by disconnected rigid chains (blue). c, f, Corresponding plastic metamaterials at different stages of compression. Arrows of red intensity indicate horizontal displacement. d, g, Corresponding vertical force and average horizontal displacement, $u_{1}$ of each floppy chain vs. global compressing stroke $u/H$. $B$, $M$, and $T$ denote the bottom, middle, and top floppy chains. The sign of the slope ($S$) of force vs compression is detected (dashed lines).
@@
Sequential Yield Buckling – We separate the excitation of the identical modes by exploiting yield buckling [26], which is rooted in the symbiotic occurrence of plastic yielding and buckling. To combine plasticity with the floppy modes, we 3D print such a metamaterial with an elastoplastic material (see Supplementary Material). Under quasistatic compression, the bottom floppy mode buckles first, followed by a negative stiffness, until all the diamonds in the first buckled layer reach self-contact (Fig. 3c top). At this point, the metamaterial stiffens until the critical buckling load of the single floppy mode is reached again (Fig. 3c bottom). The sequential deformation of the two floppy modes allows going past the isolated inner objects and the curved boundary, and enables stable energy absorption (Fig. 3d). Such sequential yield buckling of the identical floppy modes can also be designed in a large-scale metamaterial with a higher number of buckling sequences and can be generalized to other elastoplastic materials (see Appendix). However, since the geometry of each floppy mode is identical, the buckling order of the modes is determined by imperfections and boundary conditions.
We further demonstrate that the length of floppy modes can harness the buckling order in metamaterials. We design three roughly horizontal floppy chains of different lengths and separate them by rigid layers (Fig. 3e). Under quasistatic compression, the straight mode of shortest length buckles first, followed by a negative stiffness, until all the diamonds in the first buckled layer reach self-contact (Fig. 3f top). At this point, the metamaterial stiffens until the critical buckling load of the second-length mode is reached (Fig. 3f middle). Such sequential buckling repeats until the last mode with the longest length buckles and reaches contact (Fig. 3f bottom). Due to this sequential buckling, the force curve exhibits a wiggly increasing plateau with tunable local maxima (Fig. 3g). In contrast, in a metamaterial of the same geometry, but printed from an elastic material, the three floppy modes buckle almost simultaneously because of their close buckling loads and the positive stiffness at the onset of elastic buckling (see Appendix). Thus, our designed floppy modes together with plasticity bring a new approach to design the way the metamaterial collapses under compression and to tailor its nonlinear force displacement curve. Our design tool could complement optimization tools, which so far have relied on a limited set of metamaterial topologies [60, 61].
@@
Matrix-Vector Multiplication – Beyond the design of nonlinear responses, our approach can also be used to compute using mechanics; Computing in materia is an emerging direction [62, 35], studied in a myriad of physical platforms, such as cross-bar arrays [63], spintronic devices [64], electromagnetic [65, 66] and phononic coupled resonators [67], and microfluidics [68]. For these, the ability to perform algebraic operations is paramount, and therefore a crucial challenge is how to achieve matrix manipulations. Recent work has shown that floppy modes can be used to achieve matrix-vector multiplication in mechanical systems [38], which could be leveraged to create energy-efficient devices that perform complex tasks, such as battery-less speech detectors [69] or soft robots with embodied intelligence [70]. Yet, design approaches for mechanical matrix-vector products remain largely unexplored. Here, we demonstrate that floppy modes and frustrated loops are an efficient way to design structures that perform algebraic operations with multiple inputs and outputs.
We focus on one of the most basic matrix-vector multiplication: the multiplication of a $2\times 1$ vector by a $2\times 2$ matrix, $\begin{pmatrix}u&v\end{pmatrix}^{T}=M\begin{pmatrix}x&y\end{pmatrix}^{T}$. To do this, we construct a family of minimal metamaterials made from six triangular blocks (Fig. 4a-c). They all include either an open chain or a loop. We fix the hexagon’s six corner nodes and locally actuate the top-left and top-right edge nodes, as input displacements $x$ and $y$. We isolate the left and right edge nodes with two $T_{1}$ blocks to have only two outputs, $u$ and $v$, at the bottom-left and bottom-right edges. The matrix $M$ hence encodes the deformation of the chain or loop.
The ideal models of rigid bonds, realized above using LEGO®, are either purely floppy—for open chains and even loops—or strictly rigid—for odd loops. Hence, the entries of $M$ would be limited to either $\pm 1$ or identically zero, respectively. Luckily however, 3D-printed metamaterials are made of hinges that shear and extend [71, 72, 49, 43]. These elastic distortions induce a decay in displacement by some factor $\alpha<1$ from one node to the next and, in turn, non-trivial output displacements (Fig. 4a-c and Video 4). These displacements add up to form the matrix $M$ with coefficients that are rational functions of $\alpha$ which are set by the length and the topology of the loop; One can determine the sign of the interaction by patterning with an odd number of antiferromagnetic interactions—which flips the displacement’s direction. One can determine the value of the coefficient by choosing the length of the chain (see Appendix). Hence, the combinatorial design of the loops yields large design freedom for the matrix $M$.
We conclude by testing the matrix-vector product multiplication experimentally. We actuate each of the inputs $x$ and $y$ separately and observe linear behavior of each of the outputs $u$ and $v$, as expected (Fig. 4d-f). Remarkably, our measurements are well captured by our analytical predictions, both in the case of single (Fig. 4d-f), or combined input displacements (Fig. 4g-i). Our approach can readily be generalized to larger matrices (see Supplementary Material) and further design freedom can be gained by considering 3D stacks of combinatorial metamaterials, each made of multiple minimal hexagons (see Supplementary Material). Thus, our combinatorial strategy is a powerful way to design on demand matrix-vector multiplication.
@@
Figure 4: Matrix-vector multiplication. a-c, Experimental setup of open chain (a), even loop (b), and odd loop (c), with the six corner nodes fixed in place. Arrows indicate the decay in displacement along the chain or loop from a single input $x$. Solid and dashed arrows for the odd loop indicate the decay along the two paths. d-f Experimental results (symbols) and theoretical predictions (dashed lines) of output under a single input $x$ with $\alpha=0.92$ for the open chain and even loop, and $\alpha=0.8$ for the odd loop. Colors and markers correspond to the outputs marked in panels a-c. g-i, Experimental measurements (symbols) and theoretical predictions (dashed lines) under simultaneous actuation of the two inputs $x$ and $y$.
@@
We introduced a combinatorial approach using unimodal and bimodal triangular building blocks to create metamaterials with arbitrary shapes of multiple floppy modes and frustrated loops. We demonstrated that curved floppy modes together with yield buckling can be used to achieve sequential deformations, and that frustrated loops can be exploited to achieve matrix-vector multiplication. Our findings enrich the toolbox of sequential metamaterial design and lead to various applications in shock and vibration damping, shape changing, and mechanical computing. The geometric framework we have introduced opens exciting questions: how to leverage geometric nonlinearities to create even more advanced functions such as non-reciprocal computing [73] and memory devices [52]? How to harness material nonlinearities such as viscoelasticity, plasticity, and fatigue to endow mechanical computing with memory [74, 75]?
@@
Acknowledgments – We thank Chaviva Sirote-Katz, Daniela Kraft, Dor Shohat, Julio Melio, Marc Serra-Garcia, and Yuan Zhou for inspiring discussions and Bat-El Pinchasik and Yaara Shokef for technical assistance. T.A.S. and Y.S. thank the Institute of Physics at the University of Amsterdam for its hospitality. T.A.S. acknowledges funding from the Israeli Ministry of Energy and Infrastructure’s scholarship program for undergraduate students in the field of energy. W.L. and C.C. acknowledge funding from the European Research Council under grant agreement 852587 and the Netherlands Organisation for Scientific Research under grant agreement NWO TTW 17883. C.C. acknowledges funding from the Netherlands Organisation for Scientific Research under grant agreement NWO under grant agreement VI.Vidi 2131313. This research was supported in part by the Israel Science Foundation Grant No. 1899/20.
@@
Code and Data Availability – The supplementary videos as well as all the data and computer codes supporting this work are available at https://zenodo.org/records/17233065.